# Grounded 3D-LLM with Referent Tokens

### v1

@@ Appendix B Implementation Details
Contrastive language-scene pre-training.
The pre-training is conducted on Grounded Scene Caption, ScanNet-200 detection data, ScanRefer, and Multi3DRef data unless otherwise specified. Following the prior detector settings [62], we initialize 150 queries via farthest point sampling (FPS) from the entire scene. The temperature for all contrastive learning activities is set at 0.1. CLASP is initialized using the pre-trained 3D detector Mask3D [62], incorporating similar mask losses as follows:
$\mathcal{L}_{mask}=\lambda_{SF}\mathcal{L}_{SF}+\lambda_{dice}\mathcal{L}_{dice},$ (4)
where $\mathcal{L}_{SF}$ represents the sigmoid focal loss, and $\mathcal{L}_{dice}$ denotes the Dice loss for segmentation. We set $\lambda_{BCE}=5$ and $\lambda_{dice}=2$. The overall CLASP loss is defined as:
where $S_{mask}$ represents the similarity matrix between instances and points, and $S_{text}$ is the similarity matrix between text phrases and instances. $T_{mask}$ and $T_{text}$ are the corresponding ground-truth mappings. Note that $T_{mask}$ is computed using Hungarian matching. $\mathcal{L}_{cls}$ represents the sigmoid focal loss. We omit the background class because it does not correspond to a well-defined text prompt. Further implementation details can be found in Mask3D [62].
Multi-task instruction fine-tuning.
During instruction fine-tuning, the LLM is frozen, except for the projection and language heads. We utilize LoRA [32] to fine-tune the LLM. Grounded 3D-LLM is initialized with a pre-trained CLASP network. To minimize token costs, only 100 queries are utilized. The instruction fine-tuning dataset includes data conversation using various task-specific instruction templates as detailed in Sec. G. The converted datasets comprises ScanRefer, Multi3DRef, ScanNet-200 detection, along with Scan2Cap and ScanQA dataset. During training, a visual query is considered a positive match only if its mask IoU with the ground-truth object is greater than $0.3$. An additional projection layer is employed to align these features with the language embedding space of the LLM. The overall language loss is used as follows:
$\mathcal{L}_{LLM}=\mathcal{L}_{\text{language}}+\mathcal{L}_{ref}(S_{ref},T_{ref}),$ (6)
where $L_{\text{language}}$ is the default cross-entropy loss as LLAMA [64] and $L_{ref}$ computes the sigmoid focal loss between the similarity matrix of referent tokens $S_{ref}$ and the ground-truth referent-instance mapping $T_{ref}$.
Appendix C Additional Ablation Studies
Designs of cross-modal interactor. As illustrated in Table A-8, we evaluate the architectures of the scene-text interaction network, which includes cross-modal interactions using bidirectional cross attentions [49] subsequent to the BERT encoder and the prevalent Q-Former structure [43]. The results show that incorporating bidirectional cross-attention modules in cross-modal interactor enhances performance on grounding metrics and dense captioning metrics. This enhancement indicates that the module derives greater benefits from CLASP pre-training, thereby establishing it as the default cross-modal interactor.
Table 8: Comparison of cross-modal interactors. “Bi-Direct C. A.” denotes the use of bidirectional cross-attention modules (default in the cross-modal interactor) for interacting with scene queries following a BERT encoder. “Q-Former” resembles BLIP-2 [43, 17] but incorporates our scene queries.
Cross- interactor types ScanRefer Scan2Cap
Acc@0.25 Acc@0.5 BLEU-4@0.5 CIDEr@0.5
Q-Former 47.2 44.0 34.9 70.4
Bi-Direct C. A. 47.9 44.1 35.5 70.6
Diversity of instruction templates. Typical computer vision tasks such as language grounding, object detection, and dense captioning require instruction templates to better adapt to pre-trained weights of large language models. We assess the impact of diverse instruction templates on language tasks in Table A-9. Results show that diverse instruction templates consistently yield significant improvements in tasks like dense captioning and 3D QA. Notably, in template-based captioning tasks (Scan2Cap), both CIDEr@0.25 and CIDEr@0.5 see improvements exceeding $8$ points. Employing a single instruction template also leads to unstable results in repeated experiments if diverse templates are not used. Therefore, the diversity of instruction-tuning templates is crucial for sustaining language capabilities in visual tasks.
Table 9: Comparison with and without using diverse instruction templates.
Diverse Templates Scan2Cap ScanQA
B-4@0.25 C@0.25 B-4@0.5 C@0.5 B-4 C
31.5 65.8 29.2 62.1 11.2 70.8
37.2 74.4 35.5 70.6 13.4 72.7
Effects of LLM model sizes. As shown in Table A-10, we compare the performance of GPT-2, Tiny-Vicuna-1B, Vicuna-7B, and Vicuna-13B. Smaller models like GPT-2 struggle with soft referent token learning. Conversely, Grounded 3D-LLM models, ranging from 1B to 13B parameters, greatly outperform GPT-2 and show marginal improvements in dense captioning tasks, yet they experience a slight decline in 3D visual QA performance. Notably, even when only the projection layer is trainable (without LoRA), Grounded 3D-LLM (Vicuna-7B) still achieves comparable performance. Overall, these models do not significantly excel in language tasks, suggesting that models up to Vicuna-1B are adequate for existing grounding benchmarks, indicating that the current task scope may not fully challenge larger-scale LLMs.
Table 10: Comparison of different LLMs. For GPT-2, we train the overall backbones with a smaller language head. In models without LoRA, only the projection layers are trained.
LLM LoRA Scan2Cap ScanQA Trainable Params
B-4@0.25 C@0.25 B-4@0.5 C@0.5 B-4 C
GPT-2 25.1 36.7 24.2 35.2 4.91 41.9 169 MB
Tiny-Vicuna-1B 37.0 74.4 35.5 70.6 13.4 72.7 157 MB
Vicuna-7B 38.2 75.6 36.0 71.2 12.3 71.8 364 MB
Vicuna-7B 38.1 74.3 35.1 71.0 11.8 70.6 330 MB
Vicuna-13B 39.5 75.4 36.1 69.2 12.1 71.6 487 MB
Appendix D Result Visualizations
@@
Result visualization. We visualize the results in Fig. A-4 for a single model Grounded 3D-LLM that solves all downstream tasks, as opposed to per-task fine-tuning on downstream tasks[88, 31, 11].
@@
Failure cases. As depicted in Fig. A-5, the primary reasons for our model’s failures in the grounding task include missing grounding objects (a, c), excessive grounding of semantically incorrect objects (b), and errors in detection/segmentation (c).
@@
Appendix E Grounded Scene Caption Dataset
@@
Dataset language components visualization. For the grounded scene caption dataset, we visualize in Fig. A-6 the relationships between noun-adjective pairs and spatial preposition-noun pairs across all languages, along with a word cloud of all nouns, to demonstrate the diversity of our curated dataset.
@@
Dataset visualization. We provide additional visualizations of grounded scene caption examples in Fig. A-7.
@@
Appendix F Data Generation Prompts
Fig. A-8 displays the object caption prompts, including the VLM captioning prompt [70] and the subsequent condensation of the caption into a descriptive phrase using ChatGPT. These object captions are then compiled to create local scene captions, as shown in Fig. A-9, while retaining their object IDs. Finally, spatial relationships derived from program-generated code are incorporated using the prompt depicted in Fig. A-10.
Figure 8: Object caption prompt for VLM and ChatGPT (used in step 1 of data generation). The above is the prompt for CogVLM that utilizes the object labels from ScanNet-200. Below is the prompt for ChatGPT input, where “System” denotes the system prompt and “User” denotes user input.
@@ Appendix F Data Generation Prompts
Appendix G Task-specific Instruction-tuning Templates
Grounded 3D-LLM simultaneously undertakes vision tasks—such as 3D visual grounding, 3D object detection—and language tasks, like question answering. Each task demands specific, diverse templates to enable the LLM to naturally provide scene-aware answers tailored to particular dataset specifications. We show several instruction templates in the following figures.
Object detection. The detection tasks employ various question templates for each category, along with three answer templates to accommodate various detection outcomes – single, multiple, or no objects detected – as illustrated in Fig. A-11.
Figure 11: Instruction templates for object detection.
Single and multi-object grounding. The language grounding’s question template, as depicted in Fig. A-12, directs the model to identify referred objects. For the ScanRefer grounding dataset, the answer templates indicate single-object grounding. The answer templates for the Multi3DRef dataset accommodate various grounding scenarios—single, multiple, or no grounded objects.
Figure 12: Instruction templates for single- or multi- object grounding.
Question answering. For ScanQA, we append suffixes to brief outputs (Fig. A-13) to indicate the outputs with only phrases or word output as the dataset annotation guideline.
Figure 13: Instruction templates for ScanQA.
Figure 14: Instruction templates for Scan2Cap.
Dense captioning. Question templates for Scan2Cap require the model to describe object appearance followed by spatial relations, as illustrated in Fig. A-14.
Appendix H Extension to embodied dialogue and embodied planning
@@
To demonstrate potential robot applications in embodied dialogue and planning, we trained a model using a prior instruction-tuning dataset combined with a generated embodied dialogue and planning dataset. Fig A-15 and A-16 list the prompts for grounded scene caption-based dialogue and planning. The training dataset includes 11.9K dialogue examples and 4.4K planning examples. Training setups are the same as Sec.4.1.
Result visualization.
As there is no well-defined benchmark for evaluating both phrase grounding and language understanding, we visualize several results for embodied dialogue (Fig.17) and embodied planning (Fig.18). These examples show the potential of Grounded 3D-LLM’s grounding capability to enhance robot navigation and manipulation.

### v2

@@ 7 Implementation Details
8 Result Visualizations
@@
Result visualization. We visualize the results in Fig. A-7 for a single model Grounded 3D-LLM that solves all downstream tasks, as opposed to per-task fine-tuning on downstream tasks[25, 21, 20].
@@
Failure cases. As depicted in Fig. A-8, the primary reasons for our model’s failures in the grounding task include missing grounding objects (a, c), excessive grounding of semantically incorrect objects (b), and errors in detection/segmentation (c).
@@
9 Grounded Scene Caption Dataset
@@
Dataset language components visualization. For the grounded scene caption dataset, we visualize in Fig. A-9 the relationships between noun-adjective pairs and spatial preposition-noun pairs across all languages, along with a word cloud of all nouns, to demonstrate the diversity of our curated dataset.
@@
Dataset visualization. We provide additional visualizations of grounded scene caption examples in Fig. A-10.
@@
Data generation prompts Fig. A-11 displays the object caption prompts, including the VLM captioning prompt [38] and the subsequent condensation of the caption into a descriptive phrase using ChatGPT. These object captions are then compiled to create local scene captions, as shown in Fig. A-12, while retaining their object IDs. Finally, spatial relationships derived from program-generated code are incorporated using the prompt depicted in Fig. A-13.
Figure 11: Object caption prompt for VLM and ChatGPT (used in Step-1 of data generation). The above is the prompt for CogVLM that utilizes the object labels from ScanNet-200. Below is the prompt for ChatGPT input, where “System” denotes the system prompt and “User” denotes user input.
@@
10 Task-specific Instruction-following Templates
Grounded 3D-LLM simultaneously undertakes vision tasks—such as 3D visual grounding, 3D object detection—and language tasks, like question answering. Each task demands specific, diverse templates to enable the LLM to naturally provide scene-aware answers tailored to particular dataset specifications. We show several instruction-following templates in the following figures.
Object detection. The detection tasks employ various question templates for each category, along with three answer templates to accommodate various detection outcomes – single, multiple, or no objects detected – as illustrated in Fig. A-14.
Figure 14: Instruction-following templates for object detection.
Single and multi-object grounding. The language grounding’s question template, as depicted in Fig. A-15, directs the model to identify referred objects. For the ScanRefer grounding dataset, the answer templates indicate single-object grounding. The answer templates for the Multi3DRef dataset accommodate various grounding scenarios—single, multiple, or no grounded objects.
Figure 15: Instruction-following templates for single- or multi- object grounding.
3D Question answering. For ScanQA, we append suffixes to brief outputs (Fig. A-16) to indicate the outputs with only phrases or word output as the dataset annotation guideline.
Figure 16: Instruction-following templates for ScanQA.
Figure 17: Instruction-following templates for Scan2Cap.
Dense captioning. Question templates for Scan2Cap require the model to describe object appearance followed by spatial relations, as illustrated in Fig. A-17.
11 Extension to embodied dialogue and embodied planning
@@
Fig A-18 and A-19 list the prompts for grounded scene caption-based embodied dialogue and planning. The training dataset includes 11.9K dialogue examples and 4.4K planning examples.
Result visualization. As there is no well-defined benchmark for evaluating both phrase grounding and language understanding, we directly visualize the results of the jointly trained Grounded 3D-LLM for embodied dialogue (Fig. 20) and embodied planning (Fig. 21). These examples demonstrate the potential of Grounded 3D-LLM’s grounding capability to enhance robot navigation and manipulation.